# Trigonometry of partially truncated hyperbolic triangles and tetrahedra

### v1

@@ 3. Dimension three: transversals of truncated tetrahedra
Similarly to the proof of Lemma 3.1, for each $i\in\{1,2,3,4\}$ let $\mathbf{y}_{i}$ be an outward unit normal, in the sense described below Lemma 1.5, to the half-space $H_{i}$ bounded by $P_{i}$ that contains each other $P_{j}$. For any $j\neq i$, $\mathbf{y}_{i}$ and $\mathbf{y}_{j}$ are then oppositely-oriented tangent vectors to the hyperbolic geodesic intersecting $P_{i}$ and $P_{j}$ perpendicularly, so $\mathbf{y}_{i}\circ\mathbf{y}_{j}<0$ by Lemma 1.8. Furthermore, by this result the hyperbolic cosine of the distance from $P_{i}$ to $P_{j}$, which we will here denote $L_{ij}$, satisfies $L_{ij}=-\mathbf{y}_{i}\circ\mathbf{y}_{j}$ for each $i<j\in\{1,2,3,4\}$.
Since the $P_{i}$ are labeled arbitrarily, we may take $(i,j,k,l)=(1,2,3,4)$ without loss of generality. We will also prove the Lemma’s conclusion for particular parametrizations of $\tilde{\lambda}_{12}$ and $\tilde{\lambda}_{34}$ below. This will imply the general result, since for any other parametrization the resulting $D(s,t)$ will be obtained from this one by precomposing with a translation of $\mathbb{R}^{2}$ and a map of the form $(s,t)\mapsto(\pm s,\pm t)$.
@@
Let $\mathbf{v}_{1}$ and $\mathbf{v}_{3}$ be the points of intersection $\tilde{\lambda}_{12}\cap P_{1}$ and $\tilde{\lambda}_{34}\cap P_{3}$, respectively. Then $\tilde{\lambda}_{12}$ is parametrized by arclength as $\tilde{\lambda}_{12}(s)=\cosh s\,\mathbf{v}_{1}-\sinh s\,\mathbf{y}_{1}$ starting at $\mathbf{v}_{1}$ and running into $H_{1}$, since $\mathbf{y}_{1}$ is outward-pointing from $H_{1}$, and likewise $\tilde{\lambda}_{34}(t)=\cosh t\,\mathbf{v}_{3}-\sinh t\,\mathbf{y}_{3}$ starts at $\mathbf{v}_{3}$ and runs into $H_{3}$. $D(s,t)$ described above thus satisfies
@@
$=-\cosh s\cosh t\,(\mathbf{v}_{1}\circ\mathbf{v}_{3})+\cosh s\sinh t\,(\mathbf{v}_{1}\circ\mathbf{y}_{3})$
@@
We claim first that $D(s,t)\to\infty$ as $s$ and $t$ escape compact sets. To see this we write:
(6) $\frac{D(s,t)}{\cosh s\cosh t}=-(\mathbf{v}_{1}\circ\mathbf{v}_{3})+\tanh t\,(\mathbf{v}_{1}\circ\mathbf{y}_{3})+\tanh s\,(\mathbf{y}_{1}\circ\mathbf{v}_{3})-\tanh s\tanh t\,(\mathbf{y}_{1}\circ\mathbf{y}_{3}),$
and record each inner product above in terms of the distances $L_{ij}$ from $P_{i}$ to $P_{j}$ by substituting the formulas for $\mathbf{v}_{1}$ and $\mathbf{v}_{3}$ from Lemma 1.9.
@@
For large enough values of $t$, values of the ratio (6) can be made arbitrarily close to $-(\mathbf{v}_{1}\circ\mathbf{v}_{3})+(\mathbf{v}_{1}\circ\mathbf{y}_{3})+\tanh s(\mathbf{y}_{1}\circ\mathbf{v}_{3})-\tanh s(\mathbf{y}_{1}\circ\mathbf{y}_{3})$. Substituting from (7), we write this as:
$\left[\frac{L_{34}}{\sqrt{L_{34}^{2}-1}}-1\right]$ $\left[L_{13}\left(\frac{L_{12}}{\sqrt{L_{12}^{2}-1}}-\tanh s\right)+\frac{L_{23}}{\sqrt{L_{12}^{2}-1}}\right]$
$\qquad+\frac{L_{14}}{\sqrt{L_{34}^{2}-1}}\left[\frac{L_{12}}{\sqrt{L_{12}^{2}-1}}-\tanh s\right]+\frac{L_{24}}{\sqrt{(L_{12}^{2}-1)(L_{34}^{2}-1)}}$
Since $\tanh s<1$ for all $s$, this sum exceeds a fixed positive bound regardless of the value of $s$.
Now as $t$ decreases without bound, the ratio (6) approaches $-(\mathbf{v}_{1}\circ\mathbf{v}_{3})-(\mathbf{v}_{1}\circ\mathbf{y}_{3})+\tanh s(\mathbf{y}_{1}\circ\mathbf{v}_{3})+\tanh s(\mathbf{y}_{1}\circ\mathbf{y}_{3})$. This is again seen to have a positive lower bound, upon writing it as:
$\left[\frac{L_{13}L_{34}+L_{14}}{\sqrt{L_{34}^{2}-1}}+L_{13}\right]$ $\left[\frac{L_{12}}{\sqrt{L_{12}^{2}-1}}-\tanh s\right]+\frac{L_{23}L_{34}+L_{24}}{\sqrt{(L_{12}^{2}-1)(L_{34}^{2}-1)}}+\frac{L_{23}}{\sqrt{L_{12}^{2}-1}}$
We thus have a fixed positive lower bound for the ratio $D(s,t)/(\cosh s\cosh t)$, for large enough values of $|t|$. It follows that $D(s,t)\to\infty$ with $|t|$, regardless of the value of $s$. A parallel argument to the above (or again using the fact that the numbering of the $P_{i}$ is arbitrary) shows that the same is true with the roles of $s$ and $t$ reversed, proving the claim.
The claim implies that $D(s,t)$ does attain a minimum on $\mathbb{R}^{2}$, necessarily occurring at a critical point. Our next goal is to classify critical points by setting the gradient $\nabla D(s,t)=\mathbf{0}$ equal to 00 and solving for $(s,t)$. Computing partials and dividing by $\cosh s\cosh t$ yields:
$\frac{\partial D/\partial s}{\cosh s\cosh t}$ $=-\tanh s\,(\mathbf{v}_{1}\circ\mathbf{v}_{3})+\tanh s\tanh t\,(\mathbf{v}_{1}\circ\mathbf{y}_{3})+\mathbf{y}_{1}\circ\mathbf{v}_{3}-\tanh t\,(\mathbf{y}_{1}\circ\mathbf{y}_{3})$
$\frac{\partial D/\partial t}{\cosh s\cosh t}$ $=-\tanh t\,(\mathbf{v}_{1}\circ\mathbf{v}_{3})+(\mathbf{v}_{1}\circ\mathbf{y}_{3})+\tanh s\tanh t\,(\mathbf{y}_{1}\circ\mathbf{v}_{3})-\tanh s\,(\mathbf{y}_{1}\circ\mathbf{y}_{3})$
Setting these two equations equal to 00 simultaneously, then solving the first for $\tanh s$ in terms of $\tanh t$ and substituting into the second, yields the quadratic $a(\tanh t)^{2}+b\,\tanh t+a=0$ for
$a=(\mathbf{y}_{1}\circ\mathbf{y}_{3})(\mathbf{y}_{1}\circ\mathbf{v}_{3})-(\mathbf{v}_{1}\circ\mathbf{y}_{3})(\mathbf{v}_{1}\circ\mathbf{v}_{3})\ \ \mbox{and}\ \ b=(\mathbf{v}_{1}\circ\mathbf{v}_{3})^{2}+(\mathbf{v}_{1}\circ\mathbf{y}_{3})^{2}-(\mathbf{y}_{1}\circ\mathbf{v}_{3})^{2}-(\mathbf{y}_{1}\circ\mathbf{y}_{3})^{2}.$
Direct computation shows that $a<0$ and that
$b+2a$ $=(\mathbf{v}_{1}\circ\mathbf{v}_{3}-\mathbf{v}_{1}\circ\mathbf{y}_{3})^{2}-(\mathbf{y}_{1}\circ\mathbf{v}_{3}-\mathbf{y}_{1}\circ\mathbf{y}_{3})^{2}$
$=(\mathbf{v}_{1}\circ\mathbf{v}_{3}-\mathbf{v}_{1}\circ\mathbf{y}_{3}-\mathbf{y}_{1}\circ\mathbf{v}_{3}+\mathbf{y}_{1}\circ\mathbf{y}_{3})(\mathbf{v}_{1}\circ\mathbf{v}_{3}-\mathbf{v}_{1}\circ\mathbf{y}_{3}+\mathbf{y}_{1}\circ\mathbf{v}_{3}-\mathbf{y}_{1}\circ\mathbf{y}_{3})$
$>0,\ \mbox{being a product of negative numbers.}$.
(The relevant computations here are very similar to those above.) It therefore follows that the roots of the quadratic above in $\tanh t$ are positive and their product is one, with one root larger than one and one smaller. Since $\tanh t$ takes only values less than $1$, only the smaller root corresponds to a critical point. Therefore $D(s,t)$ has a unique critical point, which must correspond to its global minimum.
The minimum of $D(s,t)$ thus occurs at a point $(s_{0},t_{0})$ that is given in terms of the $L_{ij}$ by the quadratic formula. Since the values of $D(s,t)$ depend only on $s$, $t$, and the $L_{ij}$, it follows that the minimum value itself depends only on the $L_{ij}$.∎
Lemma 3.4.
Suppose $P_{1}$, $P_{2}$, $P_{3}$, $P_{4}$ are pairwise disjoint and non-parallel planes in $\mathbb{H}^{3}$ such that for each $i$, a single half-space $H_{i}$ bounded by $P_{i}$ contains $P_{j}$ for all $j\neq i$, and let $\Delta$ be the truncated tetrahedron determined by the $P_{i}$ as in Definition 3.2. For any fixed $i<j\in\{1,2,3,4\}$ and $k<l\in\{1,2,3,4\}-\{i,j\}$, the length of the transversal of $\Delta$ joining $\lambda_{ij}$ to $\lambda_{kl}$ equals the minimum distance between the geodesics $\tilde{\lambda}_{ij}$ and $\tilde{\lambda}_{kl}$ respectively containing them. It depends only on the lengths of the internal edges of $\Delta$, and written as $T(x,y;a,b,c,d)$, where $x=\cosh\ell(\lambda_{ij})$, $y=\cosh\ell(\lambda_{kl})$, and $a$, $b$, $c$, $d$ are hyperbolic cosines of the other four internal edge lengths of $\Delta$, it is symmetric in $a$, $b$, $c$ and $d$, and strictly increasing in each of them, for any fixed $x,y>1$.
Proof.
We note that $\tilde{\lambda}_{ij}$ and $\tilde{\lambda}_{kl}$ here are the same-named geodesics from Lemma 3.3. Given this, the claim that the transversal length is the minimum distance between them implies that $T(x,y;a,b,c,d)$ equals the minimum value of $D(s,t)$ from that result. The fact that it depends only on the internal edge lengths will thus follow directly from the conclusion of Lemma 3.3—since the internal edge lengths are the pairwise distances between planes—and the symmetry will follow from the fact that the labeling of the $P_{i}$, and hence also the $L_{ij}$, is arbitrary.
To prove the claim we establish first that $\tilde{\lambda}_{ij}$ intersects $\tilde{\lambda}_{kl}$, if at all, in $\lambda_{ij}\cap\lambda_{kl}$. The two components of $\tilde{\lambda}_{ij}-\lambda_{ij}$ are contained in the half-spaces bounded by $P_{i}$ and $P_{j}$ opposite $H_{i}$ and $H_{j}$, respectively; the analog is true for the components of $\tilde{\lambda}_{kl}-\lambda_{kl}$; and this implies that each component of $\tilde{\lambda}_{ij}-\lambda_{ij}$ is entirely disjoint from $\tilde{\lambda}_{kl}$. Therefore if $\tilde{\lambda}_{ij}$ intersects $\tilde{\lambda}_{kl}$, it does so in $\lambda_{ij}\cap\lambda_{kl}$. In this case the transversal length and the minimum of $D(s,t)$ both equal 00.
We now suppose that $\tilde{\lambda}_{ij}$ does not intersect $\tilde{\lambda}_{kl}$, and let $\gamma$ be the unique shortest geodesic arc from $\tilde{\lambda}_{ij}$ to $\tilde{\lambda}_{kl}$. That is, $\gamma$ joins $\tilde{\lambda}_{ij}(s_{0})$ to $\tilde{\lambda}_{kl}(t_{0})$, where the unique minimum of $D(s,t)$ occurs at $(s_{0},t_{0})$. By a standard argument, $\gamma$ intersects each of $\tilde{\lambda}_{ij}$ and $\tilde{\lambda}_{kl}$ at a right angle (if not, one could reduce its length by moving an endpoint at a non-right angle of intersection in the direction of the smaller angle). Because $\tilde{\lambda}_{ij}$ intersects each of $P_{i}$ and $P_{j}$ at right angles, any geodesic arc with an endpoint on a component of $\tilde{\lambda}_{ij}-\lambda_{ij}$ is entirely contained in the half-space bounded by $P_{i}$ or $P_{j}$ containing that component. Such an arc is therefore disjoint from $\tilde{\lambda}_{kl}$, and it follows that $\gamma$ has its endpoints on $\lambda_{ij}$ and $\lambda_{kl}$. This proves the claim.
@@
It remains to show that $T(x,y;a,b,c,d)$ is increasing in each of its last four variables. For this, as in the proof of Lemma 3.3 we take $i=1$, $j=2$, $k=3$, and $l=4$ after renumbering the $P_{i}$ if necessary. The value $T(x,y;a,b,c,d)$ is then the minimum of the function $D(s,t)$ defined there. Although it is not explicit in the notation, the function $D$ itself depends on the values $x=L_{12}$, $y=L_{34}$, $a=L_{13}$, $b=L_{14}$, $c=L_{23}$, and $d=L_{24}$. To exhibit this, we plug the values from (7) into the formula of (6), yielding:
@@
The value $d=L_{24}$ appears only once in this formula, in the numerator of the first summand above, with a positive sign. It is thus clear that for any fixed $s$ and $t$, the value of $D(s,t)$ increases with $d$ (regarding the other $L_{ij}$ as fixed). Its absolute minimum, and hence the transversal length $T(x,y;a,b,c,d)$, therefore also increases with $d$. Since $T(x,y;a,b,c,d)$ is symmetric in $a$, $b$, $c$, and $d$, the same then holds for $a$, $b$ and $c$.∎
@@
The Proposition’s proof rests on our ability to explicitly locate the critical point of the function $D(s,t)$ from Lemma 3.3, and hence explictly compute values of $T$, in highly symmetric situations.
@@
We claim that if $a=b=c=d=L$ then the minimum of $D(s,t)$ occurs at $(s_{0},t_{0})$ where $s_{0}=\ell(\lambda_{12})/2$ and $t_{0}=\ell(\lambda_{34})/2$, yielding $\cosh s_{0}=\sqrt{\frac{1}{2}(x+1)}$ and $\cosh t_{0}=\sqrt{\frac{1}{2}(y+1)}$. This can be proved by substituting directly into the formulas for $\partial D/\partial s$ and $\partial D/\partial t$ from the proof of Lemma 3.3, showing that this $(s_{0},t_{0})$ is the unique critical point of $D$. Plugging it into the formula for $D$ from the Lemma yields:

### v2

@@ 3. Dimension three: transversals of truncated tetrahedra
Similarly to the proof of Lemma 3.1, for each $i\in\{1,2,3,4\}$ let $\mathbf{y}_{i}$ be an outward unit normal, in the sense described below Lemma 1.6, to the half-space $H_{i}$ bounded by $P_{i}$ that contains each other $P_{j}$. For any $j\neq i$, $\mathbf{y}_{i}$ and $\mathbf{y}_{j}$ are then oppositely-oriented tangent vectors to the hyperbolic geodesic intersecting $P_{i}$ and $P_{j}$ perpendicularly, so $\mathbf{y}_{i}\circ\mathbf{y}_{j}<0$ by Lemma 1.9. Furthermore, by this result the hyperbolic cosine of the distance from $P_{i}$ to $P_{j}$, which we will here denote $L_{ij}$, satisfies $L_{ij}=-\mathbf{y}_{i}\circ\mathbf{y}_{j}$ for each $i<j\in\{1,2,3,4\}$. Since the $P_{i}$ are labeled arbitrarily, we may take $(i,j,k,l)=(1,2,3,4)$ without loss of generality.
@@
Let $\mathbf{v}_{1}=\tilde{\lambda}_{12}(0)$ and $\mathbf{v}_{3}=\tilde{\lambda}_{34}(0)$ be the points of intersection $\tilde{\lambda}_{12}\cap P_{1}$ and $\tilde{\lambda}_{34}\cap P_{3}$, respectively. Then $\tilde{\lambda}_{12}$ is parametrized by arclength as $\tilde{\lambda}_{12}(s)=\cosh s\,\mathbf{v}_{1}-\sinh s\,\mathbf{y}_{1}$ starting at $\mathbf{v}_{1}$ and running into $H_{1}$ using (1), since $\mathbf{y}_{1}$ is outward-pointing from $H_{1}$, and likewise $\tilde{\lambda}_{34}(t)=\cosh t\,\mathbf{v}_{3}-\sinh t\,\mathbf{y}_{3}$ starts at $\mathbf{v}_{3}$ and runs into $H_{3}$. $D(s,t)$ described above thus satisfies
@@
(7) $=-\cosh s\cosh t\,(\mathbf{v}_{1}\circ\mathbf{v}_{3})+\cosh s\sinh t\,(\mathbf{v}_{1}\circ\mathbf{y}_{3})$
@@
$=C_{++}e^{s}e^{t}+C_{+-}e^{s}e^{-t}+C_{-+}e^{-s}e^{t}+C_{--}e^{-s}e^{-t},$
where:
(8) $C_{++}$ $=\frac{1}{4}[-\mathbf{v}_{1}\circ\mathbf{v}_{3}+\mathbf{v}_{1}\circ\mathbf{y}_{3}+\mathbf{y}_{1}\circ\mathbf{v}_{3}-\mathbf{y}_{1}\circ\mathbf{y}_{3}]$
$C_{+-}$ $=\frac{1}{4}[-\mathbf{v}_{1}\circ\mathbf{v}_{3}-\mathbf{v}_{1}\circ\mathbf{y}_{3}+\mathbf{y}_{1}\circ\mathbf{v}_{3}+\mathbf{y}_{1}\circ\mathbf{y}_{3}]$
$C_{-+}$ $=\frac{1}{4}[-\mathbf{v}_{1}\circ\mathbf{v}_{3}+\mathbf{v}_{1}\circ\mathbf{y}_{3}-\mathbf{y}_{1}\circ\mathbf{v}_{3}+\mathbf{y}_{1}\circ\mathbf{y}_{3}]$
$C_{--}$ $=\frac{1}{4}[-\mathbf{v}_{1}\circ\mathbf{v}_{3}-\mathbf{v}_{1}\circ\mathbf{y}_{3}-\mathbf{y}_{1}\circ\mathbf{v}_{3}-\mathbf{y}_{1}\circ\mathbf{y}_{3}]$
We now record each inner product above in terms of the distances $L_{ij}$ from $P_{i}$ to $P_{j}$ by substituting the formulas for $\mathbf{v}_{1}$ and $\mathbf{v}_{3}$ from Lemma 1.10.
@@
We claim that $0<C_{++}<C_{+-},C_{-+}<C_{--}$. We first consider $C_{++}$. Plugging in for the terms of (8) and rearranging yields:
$\frac{1}{4}$ $\left[L_{13}\left(\frac{L_{12}L_{34}}{\sqrt{(L_{12}^{2}-1)(L_{34}^{2}-1)}}-\frac{L_{12}}{\sqrt{L_{12}^{2}-1}}-\frac{L_{34}}{\sqrt{L_{34}^{2}-1}}+1\right)\right.$
$\quad\left.+\frac{L_{23}}{\sqrt{L_{12}^{2}-1}}\left(\frac{L_{34}}{\sqrt{L_{34}^{2}-1}}-1\right)+\frac{L_{14}}{\sqrt{L_{34}^{2}-1}}\left(\frac{L_{12}}{\sqrt{L_{12}^{2}-1}}-1\right)+\frac{L_{24}}{\sqrt{(L_{12}^{2}-1)(L_{34}^{2}-1)}}\right]$
Here, bearing in mind that each $L_{ij}$ is greater than one since it is the hyperbolic cosine of a positive length, we conclude that each term in parentheses is positive, hence that $C_{++}>0$. Now:
$C_{-+}-C_{++}=-2\mathbf{v}_{1}\circ\mathbf{y}_{3}+2\mathbf{y}_{1}\circ\mathbf{y}_{3}=2L_{13}\left(\frac{L_{12}}{\sqrt{L_{12}^{2}-1}}-1\right)+\frac{2L_{23}}{\sqrt{L_{12}^{2}-1}}>0.$
A similar argument shows that $C_{+-}>C_{++}$. And since each of the four summands of $C_{--}$ has positive sign, it is greater than the other coefficients, proving the claim.
Given the claim, Lemma 3.3 implies assertion (1) of this result. The fact that the coefficients $C_{\pm\pm}$ depend only on the $L_{ij}$ implies the same for the coordinates of the critical point $(s_{0},t_{0})$ of $D$, and hence for the absolute minimum value $D(s_{0},t_{0})$. Lemma 3.3 also gives that $(s_{0},t_{0})\in(0,\infty)^{2}$; because $\tilde{\lambda}_{ij}$ was parametrized pointing into $H_{i}$ with $\tilde{\lambda}_{ij}(0)\in P_{i}$, the fact that $s_{0}>0$ implies that the closest point $\tilde{\lambda}_{ij}(s_{0})$ to $\tilde{\lambda}_{kl}$ lies in $H_{i}$. Likewise, $\tilde{\lambda}_{kl}(t_{0})\in H_{k}$. But by swapping the roles of $i$ and $j$, and of $k$ and $l$—which only re-parametrizes $\tilde{\lambda}_{12}$ and/or $\tilde{\lambda}_{34}$—and re-running the argument above, we find that the closest point of $\tilde{\lambda}_{ij}$ to $\tilde{\lambda}_{kl}$ also lies in $H_{j}$. Therefore it lies in the edge $\lambda_{ij}$ of $\Delta$. Likewise, the closest point of $\tilde{\lambda}_{kl}$ to $\tilde{\lambda}_{ij}$ lies in the edge $\lambda_{kl}$. Therefore the geodesic arc joining these two points is the relevant transversal of $\Delta$, as asserted in (3) above.
The transversal’s length $T(x,y;a,b,c,d)$ is therefore the arccosh of the minimum value of $D(s,t)$, so the fact that it depends only on the internal edge lengths follows directly from descriptions of $D$ and of $(s_{0},t_{0})$ in terms of the $C_{\pm\pm}$—these depend only on the internal edge lengths ie. the pairwise distances between planes. The symmetry property of $T$ follows again from the fact that swapping the indicex of $P_{i}$ with $P_{j}$, and/or of $P_{k}$ with $P_{l}$, does not change its value. Each such swap acts on the collection $\{a,b,c,d\}$ as an even involution.
@@
It remains to show that $T(x,y;a,b,c,d)$ is increasing in each of its last four variables. For this (again taking $i=1$, $j=2$, $k=3$, and $l=4$ after renumbering the $P_{i}$ if necessary), we claim that for any fixed $(s,t)$ the value $D(s,t)$—with itself depends on the values $x=L_{12}$, $y=L_{34}$, $a=L_{13}$, $b=L_{14}$, $c=L_{23}$, and $d=L_{24}$—increases with $d=L_{24}$ (taking the other $L_{ij}$ as fixed). To see this, we plug the values from (9) into the formula of (7). Dividing out by $\cosh s\cosh t$ yields:
@@
The value $d=L_{24}$ appears only once in this formula, in the numerator of the first summand above, with a positive sign. The claim is thus clear. It follows that the absolute minimum of $D(s,t)$, and hence also the transversal length $T(x,y;a,b,c,d)$, also increases with $d$. Since values of $T(x,y;a,b,c,d)$ are invariant under a transitive group action on $\{a,b,c,d\}$, the same then holds for $a$, $b$ and $c$.∎
@@
The Proposition’s proof rests on the relative convenience of explicitly writing down the critical point of the function $D(s,t)$ from Lemma 3.3, and hence explictly computing values of $T$, in this highly symmetric case.
@@
Plugging these into the formula (6) from Lemma 3.3, we find that $\cosh s_{0}=\sqrt{\frac{1}{2}(x+1)}$ and $\cosh t_{0}=\sqrt{\frac{1}{2}(y+1)}$. (This means that $s_{0}=\ell(\lambda_{12})/2$ and $t_{0}=\ell(\lambda_{34})/2$, which one would expect from considerations of symmetry.) Plugging this point into the formula for $D$ from the Lemma yields: